# Evidence of isospin-symmetry violation in high-energy collisions of atomic nuclei – Supplementary Information

### v3

@@ Abstract
Recently, the NA61/SHINE collaboration at the CERN SPS reported evidence of isospin-symmetry violation in high-energy nuclear collisions [Nature Commun. 16, 2849 (2025)]. The effect was observed in the relative yields of charged and neutral kaons and cannot be explained by known sources of isospin symmetry breaking. In this work, we provide supplementary information on the theoretical aspects of that study. We discuss the historical background and introduce the concepts of isospin transformations and symmetry. Importantly, we relate isospin symmetry to the QCD flavour symmetry, and we present both conceptual and analytical proofs demonstrating the equality of the mean multiplicities of charged and neutral kaons for an initial ensemble of colliding systems that is invariant under charge-symmetry transformation.
@@ 1 Introduction
In May 2025, the NA61/SHINE collaboration at the CERN SPS reported a significant excess of charged over neutral kaon production in high-energy nucleus-nucleus collisions [NA61SHINE:2023azp]. The mostly experimental paper included a theoretical interpretation of the results in terms of the evidence of the violation of isospin symmetry beyond known effects. This part of the paper was formulated in collaboration with a group of theoretical colleagues. This paper presents supplementary information on the interpretation section of the Nature Communications paper.
The paper is organised as follows. Section 2 briefly overviews the history of the isospin concept, whereas an introduction to the isospin transformation and symmetry is presented in Sec. 3. In Sec. 4 we recall that the isospin symmetry follows from flavour symmetry of up and down quarks. The method used by NA61/SHINE for testing the isospin symmetry by checking the charge-symmetry invariance in the ensemble of charge-uniform collisions is elaborated in Secs. 5 and 6. In particular, we present the conceptual and analytical proofs of the method. Finally, in Sec. 7, we estimate the violation of the charge-symmetry invariance by different strong-interaction processes.
@@ 2 Historical background
In 1932, Heisenberg suggested that the neutron and the proton could be seen as different manifestations of the same particle, the nucleon [Heisenberg:1932dw], explaining their similar masses and interaction properties. Wigner introduced the concept of ‘isotopic spin’, or ‘isospin’ of nucleons, following the formal similarity with the more familiar spin [Wigner:1936dx] and postulated invariance (symmetry) of strong interactions with respect to rotations in the isospin space. Today, isospin is an important element of nuclear physics and nuclei classification [2006NatPh...2..311W]. The concept of isospin was promptly extended to the other strongly interacting particles, the hadrons. In particular, Kemmer [Kemmer:1939zz] applied it to the pion triplet postulated by Yukawa in 1935 [Yukawa:1935xg]. In general, when considering the hadron mass distribution, hadrons form groups (isospin multiplets) with members having similar masses but different electric charges. This suggests that strong interactions predominantly determine the masses and strong interaction properties of hadrons, whereas electromagnetic — and, to a much smaller degree, weak — interactions give rise to only small corrections. For a historical outline, see Refs. [Kemmer:1982bj, Webb].
@@
With the continuously increasing number of new hadron species being discovered, it became obvious that they are not elementary particles of strong interactions. The quark model of hadron classification proposed by Gell-Mann [Gell-Mann:1964ewy] and Zweig [Zweig:1964jf] in 1964 started a 15-year-long term in which sub-hadronic particles, quarks and gluons were discovered, and a theory of their interactions, Quantum Chromodynamics (QCD) [Fritzsch:1973pi] was established. There are six types (flavours) of quarks: up ($u$), down ($d$), strange ($s$), charm ($c$), bottom ($b$) and top ($t$) with their QCD masses increasing by five orders of magnitude. Each quark carries a colour charge (red, green or blue), while the gluons are the gauge bosons of QCD and mediate the colour interactions.
@@
Within QCD, hadrons are viewed as confined bags of quarks and gluons. They are further classified into mesons with integer spin, typically quark-antiquark states such as pions and kaons, and baryons with semi-integer spin, typically three-quark states such as protons and neutrons. However, to date, attempts to derive quantitative QCD predictions for hadron properties and multi-hadron production in high-energy collisions have not been successful. Predictions of QCD-based and QCD-inspired models suffer from large uncertainties. For recent review see Ref. [Lukashov:2024jej]. In Sec. 4, we argue that the exception here is the isospin symmetry and, in particular, its part called charge-symmetry invariance.
@@
Within QCD, the hadron isospin symmetry is considered as a result of the flavour symmetry postulated by QCD. It means that gluons interact ‘democratically’ with all quark flavours. But the results of the interactions depend heavily on the different quark masses. However, the two of them, up and down quarks, have masses close to zero ($m_{u}=2.16\pm\ 0.07$ MeV and $m_{d}=4.70\pm 0.07$ MeV [ParticleDataGroup:2024cfk]) in comparison to the typical strong interaction scale ($\Lambda_{QCD}\approx 250$ MeV [Kneur:2011vi]). Thus, one traces the hadron isospin symmetry to the $ud$-flavour symmetry [^1] The textbooks refer to this symmetry as the quark isospin symmetry. Here to avoid a possible confusion whether the symmetry is meant at the quark or hadron level, we call it the $ud$-flavour symmetry. of QCD [Pal:2014, Donoghue:1992dd]. In the limit where the $u$ and $d$ quarks can be considered massless (the so-called chiral limit), the interactions do not depend on their chirality, leading to a larger symmetry called chiral symmetry. The latter is spontaneously broken in the QCD vacuum, generating (quasi-)Goldstone bosons, pions, kaons, and the $\eta$-meson [Donoghue:1992dd, tHooft:1986ooh].
@@
Experiments show that the isospin symmetry is well fulfilled in pion-pion and pion-nucleon elastic or quasi-elastic scatterings [Pennington:2005be, Alarcon:2011zs]. Moreover, the smallness of isospin-breaking decays (for example $\eta^{\prime}\rightarrow\pi\pi\pi$ [Gross:1979ur, Serpukhov-Brussels-AnnecyLAPP:1984xib]) is observed.
@@
The above suggests that the isospin symmetry should also be well obeyed in hadron production processes. The results of NA61/SHINE [NA61SHINE:2023azp] seem to contradict this expectation [Gazdzicki:1991ih].
@@ 3 Isospin transformation and symmetry
Importantly, isospin $I$ is a property of a multiplet of states, whereas $I_{z}$ is a property of a single state. Moreover, $I_{z}$ can be inferred from the measured hadron properties using the Nishijima–Nambu–Gell-Mann (NNG) formula [Nakano:1953zz, Gell-Mann:1956iqa], which for light hadrons reads:
@@
where $Q$, $B$ and $S$ are the electric charge, baryon number and strangeness of the assembly of measured hadrons. In general, a given value of $I_{z}$ obtained from the experimental data corresponds to a distribution of $I$ values within the range:
$I_{min}\leq I\leq I_{max}\penalty 10000\ ,$ (3)
where $I_{min}=|I_{z}|$ and $I_{max}$ is the maximum total isospin when all isospin vectors are aligned. This makes the experimental test of full isospin symmetry significantly more difficult than the tests of its $I_{z}$-related sub-symmetry, the so-called charge symmetry.
@@
Let us start by considering examples for a single hadron. Here, the measurement of the hadron mass identifies the multiplet, allowing us to attribute both $I$ and $I_{z}$ of the underlying state for each measured hadron.
@@
The pair proton ($p$) and neutron ($n$) forms an ‘isodoublet’ ($I=1/2$), where $p$ has eigenvalue $I_{z}=1/2$ and $n$ has $I_{z}=-1/2$. This is also the case for the pair $K^{+}$ and $K^{0}$. Thus, an isospin transformation on the doublet $(p,n)^{T}$ mixes proton and neutron and is obtained by setting $\hat{I}_{j}=\tau_{j}/2$, with $\tau_{j}$ being the Pauli matrices. The antinucleon isodoublet $(-\bar{p},\bar{n})^{T}$ transforms just as the nucleon one (note the minus sign [Halzen:1984mc]).
@@
    Firstly, we consider an example of two nucleons. For two protons, using Eqs. 2 and 3, one gets $I_{z}=+1$ and $I=1$ ($I_{\min}=I_{\max}=1$). Similarly, for the two–neutron state, one gets $I_{z}=-1$ and $I=1$. Thus, in both cases $I$ is uniquely defined. However, for a proton–neutron state the situation is more complicated: here $I_{z}=0$, but $I_{\min}=0$ and $I_{\max}=1$, allowing for two possible values of the total isospin, $I=0$ and $I=1$. Consequently, the $|p,n\rangle$ and $|n,p\rangle$ states are not eigenstates of total isospin, as they are linear combinations of the singlet ($I=0$) and triplet ($I=1$) components. The charge–symmetry transformation changes the $|p,p\rangle$ state into the $|n,n\rangle$ one and vice versa, and similarly maps $|p,n\rangle$ into $|n,p\rangle$. This means that an ensemble of $|p,p\rangle$ and $|n,n\rangle$ states with an equal number of each is invariant under charge symmetry.
@@
    Secondly, we discuss the case of an atomic nucleus with equal numbers of protons and neutrons [^2] In terms of conserved quantities, the condition $Z=N$ is given by $Q/B=1/2$, where $Q=Z$ and $B=A$ are the nuclei’s electric charge and baryon number, respectively., $Z=N$, $A=Z+N$. For a single nucleus one gets $I_{z}=0$ and $0\leq I\leq(Z+N)/2$. Thus, the state of two nuclei $|A,A\rangle$ has $I_{z}=0$. It is transformed into itself by the charge-symmetry transformation independently of its $I$ value. Obviously, the ensemble of these states is also charge-symmetry invariant. This property will later be used to test charge-symmetry invariance in nucleus-nucleus collisions. From phenomenology of nuclear physics follows that, for almost all nuclei of the $Z=N$ type, the total isospin is zero [Lenzi2009]. Then the $|A,A\rangle$ state has $I_{z}=I=0$. This would allow extending tests of isospin symmetry beyond charge symmetry by studying nucleus-nucleus collisions.
@@
Isospin symmetry refers to the invariance of the strong interaction between hadrons under global $SU(2)$ rotation in isospin space. The transition probability from a given initial state to a given final state equals the transition probability between these states transformed by the isospin transformation. This symmetry has many empirical consequences, for example:
@@ 4 Isospin symmetry from ud-flavour symmetry
This $SU(2)$ udud-flavour (quark isospin) symmetry is only approximate because $m_{u}\neq m_{d}$. The pair $u$ and $d$ forms a ‘isodoublet’, where the quark $u$ has eigenvalue $I_{z}=1/2$ and $d$ has $I_{z}=-1/2$ (see e.g. Ref. [Pal:2014] and the quark summary in the PDG [ParticleDataGroup:2024cfk]). Thus, the $ud$-flavour transformation on the doublet $(u,d)^{T}$ mixes both quarks and is obtained by setting $\hat{I}_{j}=\tau_{j}/2$, with $\tau_{j}$ being the Pauli matrices. The antiquark isodoublet $(-\bar{d},\bar{u})^{T}$ transforms just as the quark one (note the minus sign [Halzen:1984mc]). Since hadrons are colour-singlet bound states of quarks and gluons, they inherit this symmetry: the spectrum of hadrons is arranged into isospin multiplets such as the nucleons ($p,n$) or the pions ($\pi^{+},\pi^{0},\pi^{-}$).
The fact that the $ud$-flavour transformation translates into the one for isodoublets can be seen by using, for example, microscopic currents used in lattice calculations [Dudek:2013yja], Bethe-Salpeter approaches [Maris:2003vk] and effective models of QCD [Hatsuda:1994pi, Giacosa:2024epf].
@@
Hence, in the exact $ud$-flavour symmetry limit $m_{u}=m_{d}$, the divergences vanish and the currents are conserved. Turning back to hadrons, isospin symmetry implies that only transitions preserving the total isospin $I$ and its third component $I_{z}$ are allowed. This property is known as isospin conservation.
@@
The QCD $u$ and $d$ quark masses extracted from the experimental data [^3] The QCD quark masses are not measured directly but are obtained by fitting experimental data on hadron properties. Light-quark masses are fitted within lattice QCD and chiral perturbation theory, while heavy-quark masses come from perturbative QCD analyses of quarkonium and collider data. They are always quoted in a particular renormalisation scheme. are different but small. They are much smaller than $\Lambda_{QCD}$. Thus, one argues that $ud$-flavour symmetry is well obeyed. Yet, a closer inspection of the link between the approximate $ud$-flavour and isospin symmetries shows that other features of QCD are also needed to obey isospin symmetry approximately. We discuss them below.
@@
    Neglecting at first the Goldstone bosons of QCD (pions, kaons,…), the masses of the light hadrons can be successfully described by the constituent quark model approach with $m_{u}^{\ast}\simeq m_{d}^{\ast}\sim\Lambda_{QCD}\sim 250\text{ MeV}$, see Refs. [Godfrey:1985xj, Lucha:1991vn]. For instance, the masses of conventional mesons, such as the $\rho(770)$ meson, can be described as bound states of a constituent quark and a constituent antiquark: their mass is proportional to the sum of the constituent quark masses that build it. As shown by microscopic models, such as the Nambu Jona-Lasinio one [Hatsuda:1994pi] and Bethe-Salpter approaches [Fischer:2006ub], the difference of the constituent masses is proportional to the difference of QCD masses, $m_{d}^{\ast}-m_{u}^{\ast}\simeq m_{d}-m_{u}$, thus suppressed. For example, the physical $\rho(770)$-meson masses lead to the mass difference $(m_{\rho^{+}}-m_{\rho^{0}})=-0.7\pm 0.8$ MeV. Within this context, one can easily understand that $m_{u,d}\ll\Lambda_{QCD}$ implies that isospin symmetry is well fulfilled. The symmetry is also well respected in decays and scattering properties; see Refs. [Giacosa:2024epf, Kovacs:2024cdb].
@@
    where $c_{K}$ is an appropriate constant that depends on the quark condensates, e.g. [Giacosa:2024epf]. Thus, the similarity of $K^{+}$ and $K^{0}$ masses (and similarly $K^{-}$ and $\bar{K}^{0})$ is reflected in the QCD masses of up, down and strange quarks. In particular, the strange quark mass has to be large enough, $m_{s}=93.5\pm 0.8$ MeV [ParticleDataGroup:2024cfk]. For a cross-check, one may evaluate the ratio
@@
    which compares well with the experimental value $0.99209\pm 0.00004$ (the difference is likely due to residual electromagnetic interactions). Using the statistical approach, one can estimate the ratio of mean multiplicities of charged and neutral kaons caused by the mass difference as
@@
    where $T=150$ MeV is the temperature parameter of the kaon production process. This estimate assumes direct kaon production. The detailed estimate of the effect, including resonance decays, is presented in Sec. 7. It is about several $\%$.
@@
    A special remark about the neutral members of the pseudoscalar nonet is needed. The neutral pion carries a mass $m_{\pi^{0}}^{2}\propto(m_{u}+m_{d})$ just as the charged ones. The isoscalar members $\eta_{N}=(u\bar{u}+d\bar{d})/\sqrt{2}$ and $\eta_{s}=$ $s\bar{s}$ develop a mass $m_{\eta_{N}}^{2}\propto(m_{u}+m_{d})+2c_{A}$ and $m_{\eta_{S}}^{2}\propto m_{s}+c_{A}$, as well as a mixing proportional to $\sqrt{2}c_{A}$. Here $c_{A}$ is a constant proportional to the $U(1)_{A}$ chiral (or axial) anomaly. The latter refers to a specific $U(1)_{A}$-chiral-phase symmetry of quarks that is a symmetry of the classical QCD, which is broken by gluonic quantum fluctuations [tHooft:1986ooh, tHooft:1976rip]. Since the coefficient $c_{A}$ is large (in fact dominant compared to the contributions of the QCD quark masses), the masses of the isoscalar mesons $\eta_{N}$ and $\eta_{S}$ become much larger than that of the neutral pion $\pi^{0}$. The corresponding states effectively decouple from it. Instead, $\eta_{N}$ and $\eta_{S}$ mix to form the physical mesons $\eta(547)$ and $\eta^{\prime}(958)$. The latter is particularly massive and cannot be regarded as a Goldstone boson. The neutral pion $\pi^{0}$ is almost degenerate with the charged states $\pi^{\pm}$, forming an isospin triplet. As noted in Refs. [Gross:1979ur, Pisarski:1983ms], the situation would have been different if the chiral anomaly had been small. In this case, the $\pi^{0}$ and the $\eta_{N}$ states would be almost mass degenerate. A residual $\pi^{0}$-$\eta_{N}$ small mixing term, proportional to $(m_{d}-m_{u})$, would generate a large mass splitting. The lighter state would have a mass of about $77$ MeV. Thus, pions would not form an isospin triplet of almost degenerate masses. Colloquially, we may say that the large chiral anomaly protects isospin symmetry at the level of the pions [^4] This is not the case for other nonets. Even if the chiral anomaly is non-zero, it is never the dominating component as for the $\eta^{\prime}$-meson [Giacosa:2023fdz, Giacosa:2017pos].. Importantly, the charge symmetry is not affected by the chiral anomaly.
@@
    The magnitude of isospin symmetry violation cannot be calculated exactly in QCD. For mean kaon multiplicities, it can be estimated to be about several per cent; see Sec. 7 for details.

### v4

# arXiv:2312.07176v4[nucl-th] 02 Mar 2026

@@ Abstract
Recently, the NA61/SHINE collaboration at the CERN SPS reported evidence of isospin-symmetry violation in high-energy nuclear collisions [Nature Commun. 16, 2849 (2025)]. The effect was observed in the relative yields of charged and neutral kaons and cannot be explained by known sources of isospin symmetry breaking. In this work, we extend the theoretical and phenomenological aspects of that study. We discuss the historical background and introduce the concepts of isospin transformations and symmetry. Importantly, we relate isospin symmetry to the QCD flavour symmetry, and we present both conceptual and analytical proofs demonstrating the equality of the mean multiplicities of charged and neutral kaons for an initial ensemble of colliding systems that is invariant under charge-symmetry transformation.
@@ 1 Introduction
In May 2025, the NA61/SHINE collaboration at the CERN SPS published a significant excess of charged over neutral kaon production in high-energy nucleus-nucleus collisions [6]. For completeness, we recall the key plot of the paper (see Fig. 1), which shows the ratio of charged to neutral kaons as a function of collision energy. The NA61/SHINE results are presented alongside a compilation of worldwide data.
The largely experimental paper included a theoretical interpretation of the results in terms of evidence for the violation of isospin symmetry beyond known effects. The latter part of the paper was formulated in collaboration with a group of theoretical colleagues. This paper extends the theoretical and phenomenological aspects of the experimental discovery reported in Nature Communications.
The paper is organised as follows. Section 2 briefly overviews the history of the isospin concept, whereas an introduction to the isospin transformation and symmetry is presented in Sec. 3. In Sec. 4 we recall that the isospin symmetry follows from the flavour symmetry of up and down quarks. The method used by NA61/SHINE for testing the isospin symmetry by checking the charge-symmetry invariance in the ensemble of charge-uniform collisions is elaborated in Secs. 5 and 6. In particular, we present the conceptual and analytical proofs of the method. In Sec. 7, we estimate the charge-symmetry violation (CSV) within the Hadron Resonance Gas (HRG) model, and we discuss the formally analogous case of $D$ meson production. Finally, in Sec. 8 we present closing remarks and some selected outlook.
Figure 1: Ratio $R_{K}$ of charged to neutral $K$ meson yields in nucleus-nucleus collisions as a function of collision energy. The lines show model predictions; see below for explanations. Figure adapted from Ref. [6] with new preliminary Ar+Sc 8.8 GeV point added [8].
@@ 2 Historical background
In 1932, Heisenberg suggested that the neutron and the proton could be seen as different manifestations of the same particle, the nucleon [35], explaining their similar masses and interaction properties. Wigner introduced the concept of ‘isotopic spin’, or ‘isospin’ of nucleons, following the formal similarity with the more familiar spin [57] and postulated invariance (symmetry) of strong interactions with respect to rotations in the isospin space. Today, isospin is an important element of nuclear physics and nuclei classification [55]. The concept of isospin was promptly extended to the other strongly interacting particles, the hadrons. In particular, Kemmer [36] applied it to the pion triplet postulated by Yukawa in 1935 [61]. In general, when considering the hadron mass distribution, hadrons form groups (isospin multiplets) with members having similar masses but different electric charges. This suggests that strong interactions predominantly determine the masses and strong interaction properties of hadrons, whereas electromagnetic — and, to a much smaller degree, weak — interactions give rise to only small corrections. For a historical outline, see Refs. [37, 56].
@@
With the continuous discovery of new hadron species, it has become clear that they are not elementary particles of strong interactions. The quark model of hadron classification proposed by Gell-Mann [24] and Zweig [62] in 1964 started a 15-year-long term in which sub-hadronic particles, quarks and gluons were discovered, and a theory of their interactions, Quantum Chromodynamics (QCD) [20] was established. There are six types (flavours) of quarks: up ($u$), down ($d$), strange ($s$), charm ($c$), bottom ($b$), and top ($t$) with their QCD masses increasing by five orders of magnitude. Each quark carries a colour charge (red, green, or blue), while the gluons are the gauge bosons of QCD and mediate the colour interactions.
@@
Within QCD, hadrons are viewed as confined bags of quarks and gluons. They are further classified into mesons with integer spin, typically quark-antiquark states such as pions and kaons, and baryons with semi-integer spin, typically three-quark states such as protons and neutrons. However, to date, attempts to derive quantitative QCD predictions for hadron properties and multi-hadron production in high-energy collisions have not been successful. Predictions of QCD-based and QCD-inspired models suffer from large uncertainties. For recent review see Ref. [42]. In Sec. 4, we argue that the exception here is the isospin symmetry and, in particular, its part called charge-symmetry invariance.
@@
Within QCD, the hadron isospin symmetry is considered as a result of the flavour symmetry postulated by QCD. It means that gluons interact ‘democratically’ with all quark flavours. But the results of the interactions depend heavily on the different quark masses. However, the two of them, up and down quarks, have masses close to zero ($m_{u}=2.16\pm\ 0.07$ MeV and $m_{d}=4.70\pm 0.07$ MeV [46]) in comparison to the typical strong interaction scale ($\Lambda_{QCD}\approx 250$ MeV [38]). Thus, one traces the hadron isospin symmetry to the $ud$-flavour symmetry [^1] The textbooks refer to this symmetry as the quark isospin symmetry. Here to avoid a possible confusion whether the symmetry is meant at the quark or hadron level, we call it the $ud$-flavour symmetry. of QCD [47, 14]. In the limit where the $u$ and $d$ quarks can be considered massless (the so-called chiral limit), the interactions do not depend on their chirality, leading to a larger symmetry called chiral symmetry. The latter is spontaneously broken in the QCD vacuum, generating (quasi-)Goldstone bosons, pions, kaons, and the $\eta$-meson [14, 1].
@@
Experiments show that the isospin symmetry is well fulfilled in pion-pion and pion-nucleon elastic or quasi-elastic scatterings [48, 7]. Moreover, the smallness of isospin-breaking decays (for example $\eta^{\prime}\rightarrow\pi\pi\pi$ [32, 11]) is observed.
@@
The above suggests that the isospin symmetry should also be well obeyed in hadron production processes. The results of NA61/SHINE [6] seem to contradict this expectation [22].
@@ 3 Isospin transformation and symmetry
Importantly, while isospin $I$ refers to a multiplet of states, $I_{z}$ is a property of a single state within this multiplet. Moreover, $I_{z}$ can be inferred from the measured hadron properties using the Nishijima–Nambu–Gell-Mann (NNG) formula [45, 23], which for light hadrons reads:
@@
where $Q$, $B$, and $S$ are the electric charge, baryon number, and strangeness of the assembly of measured hadrons. In general, a given value of $I_{z}$ obtained from the experimental data corresponds to a distribution of $I$ values within the range:
$I_{\min}\leq I\leq I_{\max}\penalty 10000\ ,$ (3)
where $I_{\min}=|I_{z}|$ and $I_{\max}$ is the maximum total isospin when all isospin vectors are aligned. This makes the experimental test of full isospin symmetry significantly more difficult than the tests of its $I_{z}$-related sub-symmetry, the so-called charge symmetry.
@@
Let us start by considering examples for a single hadron. Here, the measurement of hadron mass identifies the multiplet, allowing us to assign both $I$ and $I_{z}$ to the underlying state of each measured hadron.
@@
The pair proton ($p$) and neutron ($n$) forms an ‘isodoublet’ ($I=1/2$), where $p$ has eigenvalue $I_{z}=1/2$ and $n$ has $I_{z}=-1/2$. This also holds for the pair $K^{+}$ and $K^{0}$. Thus, an isospin transformation on the doublet $(p,n)^{T}$ mixes proton and neutron and is obtained by setting $\hat{I}_{j}=\tau_{j}/2$, with $\tau_{j}$ being the Pauli matrices. The antinucleon isodoublet $(-\bar{p},\bar{n})^{T}$ transforms just as the nucleon one (note the minus sign [33]).
@@
    Firstly, we consider an example of two nucleons. For two protons, using Eqs. 2 and 3, one gets $I_{z}=+1$ and $I=1$ ($I_{\min}=I_{\max}=1$). Similarly, for the two–neutron state, one gets $I_{z}=-1$ and $I=1$. Thus, in both cases $I$ is uniquely defined. However, for a proton-neutron state, the situation is more complicated: here $I_{z}=0$, but $I_{\min}=0$ and $I_{\max}=1$, allowing for two possible values of the total isospin, $I=0$ and $I=1$. Consequently, the $|p,n\rangle$ and $|n,p\rangle$ states are not eigenstates of total isospin, as they are linear combinations of the singlet ($I=0$) and triplet ($I=1$) components. The charge–symmetry transformation changes the $|p,p\rangle$ state into the $|n,n\rangle$ one and vice versa, and similarly maps $|p,n\rangle$ into $|n,p\rangle$. This means that an ensemble of $|p,p\rangle$ and $|n,n\rangle$ states with an equal number of each is invariant under charge symmetry.
@@
    Secondly, we discuss the case of an atomic nucleus with equal numbers of protons and neutrons [^2] In terms of conserved quantities, the condition $Z=N$ is given by $Q/B=1/2$, where $Q=Z$ and $B=A$ are the nuclei’s electric charge and baryon number, respectively., $Z=N$, $A=Z+N$. For a single nucleus one gets $I_{z}=0$ and $0\leq I\leq(Z+N)/2$. Thus, the state of two nuclei $|A,A\rangle$ has $I_{z}=0$. It is transformed into itself by the charge-symmetry transformation independently of its $I$ value in the phenomenology of nuclear physics; these states are also charge-symmetry invariant. This property will later be used to test charge-symmetry invariance in nucleus-nucleus collisions. From phenomenology of nuclear physics follows that, for almost all nuclei of the $Z=N$ type, the total isospin is zero [40]. Then the $|A,A\rangle$ state has $I_{z}=I=0$. This would allow the extension of tests of isospin symmetry beyond charge symmetry by studying nucleus-nucleus collisions.
@@
Isospin symmetry refers to the invariance of the strong interaction between hadrons under global $SU(2)$ rotation in isospin space. The transition probability from a given initial state to a given final state equals the transition probability between the states obtained by applying the isospin transformation to these states. This symmetry has many empirical consequences, for example:
@@ 4 Isospin symmetry from ud-flavour symmetry
This $SU(2)$ udud-flavour (quark isospin) symmetry is only approximate because $m_{u}\neq m_{d}$. The pair $u$ and $d$ forms an ‘isodoublet’, where the quark $u$ has eigenvalue $I_{z}=1/2$ and $d$ has $I_{z}=-1/2$ (see e.g. Ref. [47] and the quark summary in the PDG [46]). Thus, the $ud$-flavour transformation on the doublet $(u,d)^{T}$ mixes both quarks and is obtained by setting $\hat{I}_{j}=\tau_{j}/2$, with $\tau_{j}$ being the Pauli matrices. The antiquark isodoublet $(-\bar{d},\bar{u})^{T}$ transforms just as the quark one (note the minus sign [33]). Since hadrons are colour-singlet bound states of quarks and gluons, they inherit this symmetry: the spectrum of hadrons is arranged into isospin multiplets such as the nucleons ($p,n$) or the pions ($\pi^{+},\pi^{0},\pi^{-}$).
The fact that the $ud$-flavour transformation translates into the one for isodoublets can be seen by using, for example, microscopic currents used in lattice calculations [16], Bethe-Salpeter approaches [44] and effective models of QCD [34, 27].
@@
Hence, in the exact $ud$-flavour symmetry limit $m_{u}=m_{d}$, the divergences vanish and the currents are conserved. Turning back to hadrons, isospin symmetry implies that only transitions preserving the total isospin $I$ and its third component $I_{z}$ are allowed. This property is the already mentioned isospin conservation.
@@
The QCD $u$ and $d$ quark masses extracted from the experimental data [^3] The QCD ‘current’ quark masses $m_{u}=2.16\pm 0.07$ MeV and $m_{d}=4.70\pm 0.07$ are not measured directly but are obtained by fitting experimental data on hadron properties [46]. Light-quark masses are fitted within lattice QCD and chiral perturbation theory, while heavy-quark masses come from perturbative QCD analyses of quarkonium and collider data. They are always quoted in a particular renormalisation scheme. are different but small. They are much smaller than $\Lambda_{QCD}$. Thus, one argues that $ud$-flavour symmetry is well obeyed. Yet, a closer inspection of the link between the approximate $ud$-flavour and isospin symmetries shows that other features of QCD are also needed to obey isospin symmetry approximately. We discuss them below.
@@
    Neglecting at first the Goldstone bosons of QCD (pions, kaons,…), the masses of the light hadrons can be successfully described by the constituent quark model approach with $m_{u}^{\ast}\simeq m_{d}^{\ast}\sim\Lambda_{QCD}\sim 250\text{ MeV}$, see Refs. [30, 41]. For instance, the masses of conventional $\bar{q}q$ mesons, such as the $\rho(770)$ meson, can be described as bound states of a constituent quark and a constituent antiquark: their mass is proportional to the sum of the constituent quark masses that build it. As shown by microscopic models, such as the Nambu Jona-Lasinio one [34] and Bethe-Salpeter approaches [18], the difference of the constituent masses is proportional to the difference of QCD masses, $m_{d}^{\ast}-m_{u}^{\ast}\simeq m_{d}-m_{u}$, thus suppressed. For example, the physical $\rho(770)$-meson masses lead to the mass difference $(m_{\rho^{+}}-m_{\rho^{0}})=-0.7\pm 0.8$ MeV. Within this context, one can easily understand that $m_{u,d}\ll\Lambda_{QCD}$ implies that isospin symmetry is well fulfilled. The symmetry is also well respected in decays and scattering properties; see Refs. [27, 39].
@@
    where $c_{K}$ is an appropriate constant that depends on the quark condensates, e.g. [27]. Thus, the similarity of $K^{+}$ and $K^{0}$ masses (and similarly $K^{-}$ and $\bar{K}^{0})$ is reflected in the QCD masses of up, down, and strange quarks. In particular, the strange quark mass has to be large enough, $m_{s}=93.5\pm 0.8$ MeV [46]. For a cross-check, one may evaluate the ratio
@@
    whose central value is close to the corresponding central experimental value $0.99209\pm 0.00004$. Yet, due to the very small errors, these values are not compatible. The reason is the small but decisive role of the electromagnetic interaction [15], which slightly increases the mass of $K^{+}$ (both constituents $u$ and $\bar{s}$ have positive charge), but slightly decreases the mass of $K^{0}$ (the constituents $d$ and $\bar{s}$ have opposite charge).
    Using the statistical approach, one can estimate the ratio of mean multiplicities of charged and neutral kaons caused by the mass difference as
@@
    where $T=150$ MeV is the temperature parameter of the kaon production process. This estimate assumes direct kaon production. The detailed estimate of the effect, including resonance decays, is presented in Sec. 7. It is about several $\%$ at $\sqrt{s}\simeq 10-20$ GeV. Note that this estimate includes both strong and electromagnetic effects that modify vacuum masses of hadrons and their branching ratios (assuming they have been measured).
@@
    A special remark about the neutral members of the pseudoscalar nonet is needed. The neutral pion carries a mass $m_{\pi^{0}}^{2}\propto(m_{u}+m_{d})$ just as the charged ones. The isoscalar members $\eta_{N}=(u\bar{u}+d\bar{d})/\sqrt{2}$ and $\eta_{s}=$ $s\bar{s}$ develop a mass $m_{\eta_{N}}^{2}\propto(m_{u}+m_{d})+2c_{A}$ and $m_{\eta_{S}}^{2}\propto m_{s}+c_{A}$, as well as a mixing proportional to $\sqrt{2}c_{A}$. Here $c_{A}$ is a constant proportional to the $U(1)_{A}$ chiral (or axial) anomaly. The latter refers to a specific $U(1)_{A}$-chiral-phase symmetry of quarks that is a symmetry of the classical QCD, which is broken by gluonic quantum fluctuations [1, 2]. Since the coefficient $c_{A}$ is large (in fact dominant compared to the contributions of the QCD quark masses), the masses of the isoscalar mesons $\eta_{N}$ and $\eta_{S}$ become much larger than that of the neutral pion $\pi^{0}$. The corresponding states effectively decouple from it. Instead, $\eta_{N}$ and $\eta_{S}$ mix to form the physical mesons $\eta(547)$ and $\eta^{\prime}(958)$. The latter is particularly massive and cannot be regarded as a Goldstone boson. The neutral pion $\pi^{0}$ is almost degenerate with the charged states $\pi^{\pm}$, forming an isospin triplet. As noted in Refs. [32, 49], the situation would have been different if the chiral anomaly had been small. In this case, the $\pi^{0}$ and the $\eta_{N}$ states would be almost mass degenerate. A residual $\pi^{0}$-$\eta_{N}$ small mixing term, proportional to $(m_{d}-m_{u})$, would generate a large mass splitting. The lighter state would have a mass of about $77$ MeV. Thus, pions would not form an isospin triplet of almost degenerate masses. Colloquially, we may say that the large chiral anomaly protects isospin symmetry at the level of the pions [^4] This is not the case for other nonets. Even if the chiral anomaly is non-zero, it is never the dominating component as for the $\eta^{\prime}$-meson [25, 26].. Importantly, the charge symmetry is not affected by the chiral anomaly.
@@
    The magnitude of isospin symmetry violation cannot be calculated exactly in QCD. For mean kaon multiplicities, the estimate is approximately several per cent; see Sec. 7 for details.